OpenStudy (anonymous):

write a function for the situation and tell if it is contionous or discrete ? A movie store sells DVDs for 10$ each,what is the cost, C, of N DVDs ?

OpenStudy (anonymous):

Hint 1: If one DVD is \(\$10\), how much is \(N\) DVDs? Try some particular values of \(N\) (2, 3, 4, for instance) if you are stuck, and try to find a pattern. Hint 2: Functions are considered discrete if they only admit integers (or natural numbers). If they can admit all real numbers and are "smooth", then they are called continuous. If your independent variable here is the number of DVDs sold \(N\), can that variable only be of the integers (or natural numbers), or the real numbers?

OpenStudy (anonymous):

A. C=10 + N;discrete B. C=10 N; discrete C. C=10 + N, continous D. C=10N ; continous

OpenStudy (anonymous):

which one is it out of those 4 ? plz nd thank you its my math final

OpenStudy (anonymous):

This is not the place to simply give out answers. Please try to work it out yourself.

OpenStudy (anonymous):

its my math final nd this test tells me wether i pass freshmen year or not it was one question its not like i posted my whole test on here